OpenStudy (anonymous):

To find the roots of a polynomial equation you should:

OpenStudy (anonymous):

are you really typing

OpenStudy (anonymous):

That's the easiest question i have ever answered :D Ok so, Here are three important theorems relating to the roots of a polynomial: (a) A polynomial of n-th degree can be factored into n linear factors. (b) A polynomial equation of degree n has exactly n roots. (c) If (x−r) is a factor of a polynomial, then x=r is a root of the associated polynomial equation Example: The cubic polynomial f(x) = 4x3 − 3x2 − 25x − 6 has degree 3 (since the highest power of x that appears is 3). This polynomial can be factored (using Scientific Notebook or similar software) and written as 4x3 − 3x2 − 25x − 6 = (x − 3)(4x + 1)(x + 2) So we see that a 3rd degree polynomial has 3 roots. The associated polynomial equation is formed by setting the polynomial equal to zero: f(x) = 4x3 − 3x2 − 25x − 6 = 0 In factored form, this is: (x−3)(4x+1)(x+2)=0 We see from the expressions in brackets and using the 3rd theorem from above, that there are 3 roots, x=3, −14, −2. In this example, all 3 roots of our polynomial equation of degree 3 are real. Since (x−3) is a factor, then x=3 is a root. Since (4x+1) is a factor, then x=−14 is a root. Since (x+2) is a factor, then x=−2 is a root.